OpenStudy (anonymous):

)The Baldwin family walked four over five of a mile in four over nine of an hour. What is their unit rate in miles per hour?

OpenStudy (anonymous):

four-ninths over four-fifths equals four over nine mile per hour four-fifths over four-ninths equals one and four-fifths miles per hour four-ninths over four-fifths equals one and four-fifths miles per hour four-fifths over four-ninths equals four over nine mile per hour

OpenStudy (arindameducationusc):

you there?

OpenStudy (anonymous):

yes.

OpenStudy (arindameducationusc):

(4/5) / (9/4)

OpenStudy (anonymous):

Do I multiply that?

OpenStudy (anonymous):

Im not good at math at all, sorry.

OpenStudy (arindameducationusc):

hmm its B or D four fifths over four ninths is sure, but for second part i am confused, wait...analysing...

OpenStudy (arindameducationusc):

9/5 miles per hour is the answer.... nut not going with options....hmmm.... can you think of something?

OpenStudy (arindameducationusc):

you there?

OpenStudy (texaschic101):

(4/5) / (4/9) = 4/5 * 9/4 = 36/20 = 1 16/20 = 1 4/5 miles per hr

OpenStudy (anonymous):

yes im here.

OpenStudy (anonymous):

And thankyou @texaschic101

OpenStudy (texaschic101):

sure thing :)